OpenStudy (karabella98):

How much distance does it take to stop a car going 30 m/s (67 mph) if the brakes can apply a force equal to one half the car’s weight?

OpenStudy (abhisar):

Hey there! Welcome to OpenStudy :-) Is mass of car given?

OpenStudy (karabella98):

Hello! No, the mass is not given.

OpenStudy (abhisar):

It's ok. We can still solve it :-)

OpenStudy (abhisar):

So, let's see what are we given, Initial speed (u) = 30 m/s Since we want to stop the car, final velocity (v) will be 0 Also the maximum breaking force can be half of the car's weight, if we assume m as the car's mass then it would be mg/2. This means retardation of the car will be g/2\(\approx\)5 Now we have to calculate the distance. We can do this by using the formula \(\sf v^2=u^2-2as\)

OpenStudy (karabella98):

I got d=900m. Is that right?

OpenStudy (abhisar):

How are you getting that? Calculate s using the formula I gave you...

OpenStudy (abhisar):

\(\sf 0^2=(30)^2-2\times 5 \times s\), find s. That would be the answer

OpenStudy (karabella98):

89m

OpenStudy (abhisar):

Yep, 90 m